A 12µF capacitor is required to store 4J of energy. Find the p.d to which the capacitor must be charged
1
Expert's answer
2020-05-07T08:41:15-0400

The energy "W" , that can be stored into a capacitor is given by the expression:

"W = \dfrac{CU^2}{2}" ,

where "C" is the capacitance and "U" is the potential diffectnce (p.d).

Express "U" from this formula and substitute numerical values:

"U = \sqrt{\dfrac{2W}{C}} = \sqrt{\dfrac{2\cdot 4J}{12\cdot 10^{-6}F}} = 816.5 V" .


Answer. 816.5 V.